A car accelerates from rest to reach a certain speed in 10 minutes. If then continues at this speed for another 10 minutes and then decelerates to rest in further 5 minutes. If the total distance covered is 17.5km. Find the speed reached.


1
Expert's answer
2022-01-25T06:49:46-0500

Given:

"l=17.5\:\rm km"

"t_1=10\:\rm min"

"t_2=10\:\rm min"

"t_3=5\:\rm min"


The average speed

"v_{\rm ave}=\frac{l}{t}=\frac{l}{t_1+t_2+t_3}"

"v_{\rm ave}=\rm\frac{17.5\: km}{(10+10+5)*1/60\: hr}=42\: km/hr"